In ∆ SPY , points A and B are midpoints of SY and SP , respectively. If = ( a+ 7) units, = (3 + 3) units,PY and m∠SYP = (6a − 2)°


1
Expert's answer
2021-02-02T04:48:14-0500

∆ SPY , points A and B are midpoints of SY and SP , respectively


"PY=(a+7)units, AB=(3+3)units,"

"m\angle SYP=(6a-2)\degree"


Then

"PY=2AB=2(3+3)=a+7"

"a=5"

"6(5)-2=28"


"m\angle SYP=28\degree"